OpenStudy (anonymous):

Which answer shows this equation in standard form? 6 – 2(x – y) = –3x + 5 A. x + 2y = –1 B. x + 2y = 1 C. x – 2y = 1 D. –5x + 2y = –1 i think its d

OpenStudy (andrewthecookie):

its D correct!

OpenStudy (anonymous):

thank you

OpenStudy (igreen):

No, that's wrong. @NunChuckBoy3101

OpenStudy (anonymous):

That answer's wrong.

OpenStudy (igreen):

\(6 - 2(x - y) = -3x + 5\) Distribute 2 into the parenthesis: \(6 - 2x + 2y = -3x + 5\) Subtract 6 to both sides: \(-2x + 2y = -3x - 1\) Add 3x to both sides, what do you get?

OpenStudy (igreen):

Can you add 3x + -2x?

OpenStudy (igreen):

It's the same thing as 3x - 2x

OpenStudy (anonymous):

ok 3x−2x+2y=−3x−1

OpenStudy (igreen):

Can you tell me what 3x - 2x is? It's just 3 - 2 and adding an 'x' at the end.

OpenStudy (anonymous):

1x

OpenStudy (anonymous):

@iGreen

OpenStudy (igreen):

Yes, and 1x is the same thing as 'x'. So we have: \(x + 2y = -1\)

OpenStudy (anonymous):

oh ok thank you

OpenStudy (igreen):

No problem.